The rate of a reaction at 25 °C is 2×+0^-3.Then the rate of reaction at 35°Cis approximately

1
Expert's answer
2022-01-28T09:30:01-0500

25=2×10-3

35=

= 35×2×10-3/25

= 2.8×10-3